# TCAN: Text-oriented Cross Attention Network for Multimodal Sentiment Analysis

### v1

@@
TCAN: Text-oriented Cross Attention Network for Multimodal Sentiment Analysis
[…]
TCAN: Text-oriented Cross Attention Network for Multimodal Sentiment Analysis. JOURNAL OF COMPUTER SCIENCE AND TECHNOLOGY
${}^{1}$chool of Information Science and Technology, Donghua University, Shanghai 201620, China
${}^{2}$, Institute of Automation, Chinese Academy of Sciences, Beijing 100190, China
${}^{3}$chool of Artificial Intelligence, University of Chinese Academy of Sciences, Beijing 100049, China
${}^{4}$niv. Grenoble Alpes, CNRS, Grenoble INP, GIPSA-lab, Grenoble 38000, France
Email: mzhou@mail.dhu.edu.cn; qweizework@gamil.com; zhouziqi2022@ia.ac.cn; kai.wang@gipsa-lab.grenoble-inp.fr
wangtong@dhu.edu.cn; yandongming@gmail.com
footnotetext:
 Regular Paper
@@
AbstractMultimodal Sentiment Analysis (MSA) endeavors to understand human sentiment by leveraging language, visual, and acoustic modalities. Despite the remarkable performance exhibited by previous MSA approaches, the presence of inherent multimodal heterogeneities poses a challenge, with the contribution of different modalities varying considerably. Past research predominantly focused on improving representation learning techniques and feature fusion strategies. However, many of these efforts overlooked the variation in semantic richness among different modalities, treating each modality uniformly. This approach may lead to underestimating the significance of strong modalities while overemphasizing the importance of weak ones. Motivated by these insights, we introduce a Text-oriented Cross-Attention Network (TCAN), emphasizing the predominant role of the text modality in MSA. Specifically, for each multimodal sample, by taking unaligned sequences of the three modalities as inputs, we initially allocate the extracted unimodal features into a visual-text and an acoustic-text pair. Subsequently, we implement self-attention on the text modality and apply text-queried cross-attention to the visual and acoustic modalities. To mitigate the influence of noise signals and redundant features, we incorporate a gated control mechanism into the framework. Additionally, we introduce unimodal joint learning to gain a deeper understanding of homogeneous emotional tendencies across diverse modalities through backpropagation. Experimental results demonstrate that TCAN consistently outperforms state-of-the-art MSA methods on two datasets (CMU-MOSI and CMU-MOSEI).
Keywordsmultimodal sentiment analysis, cross-attention, multimodal fusion, multimodal representations
1 Introduction
In recent years, the widespread use of social media and the prevalence of smartphones equipped with high-quality cameras have led to a significant increase in multimodal data sharing. This surge includes various forms such as video clips on diverse social media platforms. Such multimodal data consists of three channels: visual (image), acoustic (voice), and transcribed linguistic (text) data. Various modalities often synergize within the same data segment, offering additional cues that aid in disambiguating semantics and emotions. Recognizing the research significance of sentiment, there has been a notable surge in recent studies [1, 2, 3] focusing on the detection and examination of human sentiments. A majority of these studies leverage Multimodal Sentiment Analysis (MSA), aiming to extract sentiment information from multiple data sources. This approach seeks a more comprehensive and precise understanding of a broad spectrum of intricate human emotions, standing in contrast to traditional unimodal sentiment analysis, which typically relies on a single source of information.
@@ 1 Introduction
Fig. 1: (a) illustrates the significant sentiment analysis discrepancies using unimodality, adapted from Mult [7]. (b) shows an example of unimodal labels and multimodal labels, where the green dotted lines represent the process of backpropagation.
@@ 2 Related Work
2.1 Multimodal Sentiment Analysis
@@
Fig. 2: The framework of TCAN. Given the input multimodal data, TCAN encodes their respective shallow features $F_{m}$, where $m\in\left\{t,v,a\right\}$. In the Text-oriented Cross-attention module, TCAN exploits cross-attention and self-attention mechanisms to process text-audio pairs and text-video pairs and applies the gated mechanism to control the impact of noise and redundant information (Section 3.3). At the same time, we introduce a shared-weight encoder called the Homogeneous encoder, which is used to extract homogeneous features from unimodal for joint training (Section 3.4). Finally, we compute the final representation from the outputs of the last layer in the Text-oriented Cross-attention module and then concatenate them for MSA.
2.2 Multimodal Representation Learning
@@
2.3 Joint Learning
Joint learning is also a way of multimodal feature learning, in which we can enrich multimodal features through single-modal feature learning. Different from the methods mentioned in Section 2.2, these approaches are usually based on backward guidance. However, due to the unified multimodal annotation, it is difficult for them to directly utilize multimodal labels to participate in training. Unimodal joint learning usually has the following two methods: additional unimodal annotations and auto-generated unimodal labels. The former like Yu et al.[30] introduced fine-grained annotation of modality, which needed high labor costs. An example of the latter method is [14] which designed a unimodal label generation module based on multimodal labels and modality representations. Different from both the above methods, we propose a homogeneous encoder with shared weights. Through this encoder, we can extract unimodal homogeneous features to participate in unimodal joint training.
3 Method
@@ 3.1 Problem Statement
3.2 Unimodal Feature Extraction Module
@@
Fig. 3: The architecture of the Text-oriented Cross-attention module. The cross-attention block takes text-audio and text-video pairs as its input and the self-attention block takes text as its input. After the cross-attention block, two gates were introduced to eliminate adverse information. The memory gate decides how much proportion of the target modality’s components to be kept forwarding, and the fuse gate decides how much proportion of fused components to be injected into the target modality.
3.3 Text-oriented Cross-Attention Module
The architecture of this module is shown in Fig. 3.2. In our TCAN, text-oriented cross-attention modules are stacked $N$ times. We also evaluate the impact of stack number on the final results in the ablation experiment. The module contains a cross-attention block for text-oriented feature fusion and a self-attention block for refining text features. Additionally, a gate mechanism is incorporated to effectively integrate useful information and filter out redundancies between these two blocks. The cross-attention block takes $F_{t}^{[i]}$ and $F_{m\to t}^{[i]}$ as its inputs, where $m\in\left\{a,v\right\}$, and the superscription $i$ indicates the i-th layer. First, we perform a layer normalization (LN) on the features: $F_{m\to t}^{[i]}=\mathbf{LN}(F_{m\to t}^{[i]})$ and $F_{t}^{[i]}=\mathbf{LN}(F_{t}^{[i]})$. Then, we put them into a cross-attention (CA) block:
@@ 3.3 Text-oriented Cross-Attention Module
3.4 Unimodal Joint Learning
@@ 4 Experiments
4.1 Experimental Settings
[…]
Datasets. We evaluate TCAN on CMU-MOSI [34] and CMU-MOSEI [35] datasets. To achieve a more comprehensive comparison, the experiments are conducted under word-unaligned settings. Table 4.1 depicts the composition of these two datasets. The CMU-MOSI dataset comprises 2,199 short monologue video clips. The acoustic features are extracted at a sampling rate of 12.5 Hz, while the visual features are extracted at a sampling rate of 15 Hz. The CMU-MOSEI dataset consists of a significantly larger number of samples compared to CMU-MOSI, with a total of 22,856 movie review video clips sourced from YouTube (approximately 10 times the size of CMU-MOSI). The acoustic features in CMU-MOSEI are extracted at a sampling rate of 20 Hz, while the visual features are extracted at a sampling rate of 15 Hz. For both CMU-MOSI and CMU-MOSEI datasets, each sample is annotated with a sentiment score ranging from -3 to 3. The sentiment scores represent different levels of sentiment, including highly negative, negative, weakly negative, neutral, weakly positive, positive, and highly positive. Following previous works [14, 7], we evaluate the model performance with the following metrics: the 7-class accuracy (Acc-7), the binary accuracy (Acc-2) where zero excluded evaluation results are used, the mean absolute error (MAE), the correlation of the model’s prediction with human (Corr), and the F1 score.
@@ 4.1 Experimental Settings
Table 1: The compositions of the number of samples in the CMU-MOSI and the CMU-MOSEI dataset.
@@
Table 2: Comparison results on the CMU-MOSI. Bold is the best.
@@
Table 3: Comparison results on the CMU-MOSEI dataset. Bold is the best.
@@
4.2 Comparisons with the State-of-the-Art
@@
Table 4: Ablation study on the influence of modality on the CMU-MOSEI dataset. The unimodal results are adapted from Mult [7].
@@
Table 5: Ablation study on the influence of the center modality on the CMU-MOSI.
@@
Table 6: Ablation study on the influence of different network modules on the CMU-MOSI.
@@ 4.3 Ablation Study
We separately remove each modality to investigate the performance of our proposed model. As illustrated in Table 4.2, we can see that the performance of “T only” is remarkably better than “A only” and “V only”, which to some extent implies the importance of text for MSA. When introducing text-oriented cross-attention modules, “TV pairs” or “TA pairs”, the performance has an apparent gain. Then, combining “TV pairs” and “TA pairs” further improves the recognition performance. In addition, the performance drop is relatively large when the visual modality is removed (comparing “TV pairs” and “TA pairs” related to “TV+TA pairs”).
Fig. 4: Performance of TCAN with different hyperparameter $N$ on the CMU-MOSI. Here we use $F_{1}$ scores to show the performance.
To illustrate the dominance of the text modality, we substitute the other two modalities as the dominant modality for the experiments. As illustrated in Table 5, the acoustic- and visual-oriented models consistently experience noticeable performance degradation. These observations indicate that the text modality possesses richer semantics and less noise, resulting in better feature reinforcement of the other two modalities.
Fig. 5: Some representative examples for the visualization analysis of MSA. Red represents positive sentiment and blue represents negative sentiment.
We explore the importance of the proposed components by removing the gated mechanism and unimodal joint training separately. We can see that the gain degrades when removing one of the modules. As shown in Table 6, these observations suggest that gated mechanism and joint training modules play an important role in the result.
Finally, to investigate the impact of the number $N$ of the Text-oriented Cross-attention module on the model’s performance, we carried out experiments on the CMU-MOSI dataset using varying values for parameter $N$. The results are illustrated in Fig. 4.3. As $N$ increases, we observe a pattern in which the F1 scores initially rise and then decline, with the network achieving its optimal performance when $N=5$. We believed that the larger $N$ could result in better performance. Nevertheless, experiments indicate that an excessive number of layers may restrict the text modality’s capacity to guide the other two modalities. We should choose the appropriate network for different datasets, which is exactly what our proposed TCAN can flexibly do. When transitioning our model to a more intricate dataset, we can strategically augment the number of Text-Oriented Cross-modal Attention modules to attain optimal performance.
4.4 Visualization Analysis
@@ 5 Conclusion
Conflict of Interest
The authors declare that they have no conflict of interest.

### v2

@@
Abstract
@@
Multimodal Sentiment Analysis (MSA) endeavors to understand human sentiment by leveraging language, visual, and acoustic modalities. Despite the remarkable performance exhibited by previous MSA approaches, the presence of inherent multimodal heterogeneities poses a challenge, with the contribution of different modalities varying considerably. Past research predominantly focused on improving representation learning techniques and feature fusion strategies. However, many of these efforts overlooked the variation in semantic richness among different modalities, treating each modality uniformly. This approach may lead to underestimating the significance of strong modalities while overemphasizing the importance of weak ones. Motivated by these insights, we introduce a Text-oriented Cross-Attention Network (TCAN), emphasizing the predominant role of the text modality in MSA. Specifically, for each multimodal sample, by taking unaligned sequences of the three modalities as inputs, we initially allocate the extracted unimodal features into a visual-text and an acoustic-text pair. Subsequently, we implement self-attention on the text modality and apply text-queried cross-attention to the visual and acoustic modalities. To mitigate the influence of noise signals and redundant features, we incorporate a gated control mechanism into the framework. Additionally, we introduce unimodal joint learning to gain a deeper understanding of homogeneous emotional tendencies across diverse modalities through backpropagation. Experimental results demonstrate that TCAN consistently outperforms state-of-the-art MSA methods on two datasets (CMU-MOSI and CMU-MOSEI).
Index Terms:
Multimodal sentiment analysis, cross-attention, multimodal fusion, multimodal representations.
I Introduction
@@
Figure 1: (a) illustrates the significant sentiment analysis discrepancies using unimodality, adapted from Mult [1]. (b) shows an example of unimodal labels and multimodal labels, where the green dotted lines represent the process of backpropagation.
In recent years, the widespread use of social media and the prevalence of smartphones equipped with high-quality cameras have led to a significant increase in multimodal data sharing. This surge includes various forms, such as video clips on diverse social media platforms. Such multimodal data consists of three channels: visual (image), acoustic (voice), and transcribed linguistic (text) data. Various modalities often synergize within the same data segment, offering additional cues that aid in disambiguating semantics and emotions. Recognizing the research significance of sentiment, there has been a notable surge in recent studies [2, 3, 4, 5] focusing on the detection and examination of human sentiments. A majority of these studies leverage Multimodal Sentiment Analysis (MSA), aiming to extract sentiment information from multiple data sources [6, 7]. This approach seeks a more comprehensive and precise understanding of a broad spectrum of intricate human emotions, standing in contrast to traditional unimodal sentiment analysis, which typically relies on a single source of information.
@@ II Related Work
II-A Multimodal Sentiment Analysis
@@
II-B Multimodal Representation Learning
@@
II-C Joint Learning
Joint learning is also a way of multimodal feature learning, in which we can enrich multimodal features through single-modal feature learning. Different from the methods mentioned in Section II-B, these approaches are usually based on backward guidance. However, due to the unified multimodal annotation, it is difficult for them to directly utilize multimodal labels to participate in training. Unimodal joint learning usually has the following two methods: additional unimodal annotations and auto-generated unimodal labels. The former like Yu et al. [33] introduced fine-grained annotation of modality, which needed high labor costs. An example of the latter method is [17] which designed a unimodal label generation module based on multimodal labels and modality representations. Different from both the above methods, we propose a homogeneous encoder with shared weights. Through this encoder, we can extract unimodal homogeneous features to participate in unimodal joint training.
Figure 2: The framework of TCAN. Given the input multimodal data, TCAN encodes their respective shallow features $F_{m}$, where $m\in\left\{t,v,a\right\}$. In the Text-oriented Cross-attention module, TCAN exploits cross-attention and self-attention mechanisms to process text-audio pairs and text-video pairs and applies the gated mechanism to control the impact of noise and redundant information (Section III-C). At the same time, we introduce a shared-weight encoder called the Homogeneous encoder, which is used to extract homogeneous features from unimodal for joint training (Section III-D). Finally, we compute the final representation from the outputs of the last layer in the Text-oriented Cross-attention module and then concatenate them for MSA.
III Method
@@ III-A Problem Statement
III-B Unimodal Feature Extraction Module
@@
III-C Text-oriented Cross-Attention Module
The architecture of this module is shown in Fig. 3. In our TCAN, text-oriented cross-attention modules are stacked $N$ times. We also evaluate the impact of stack number on the final results in the ablation experiment. The module contains a cross-attention block for text-oriented feature fusion and a self-attention block for refining text features. Additionally, a gate mechanism is incorporated to effectively integrate useful information and filter out redundancies between these two blocks. The cross-attention block takes $F_{t}^{[i]}$ and $F_{m\to t}^{[i]}$ as its inputs, where $m\in\left\{a,v\right\}$, and the superscription $i$ indicates the i-th layer. First, we perform a layer normalization (LN) on the features: $F_{m\to t}^{[i]}=\mathbf{LN}(F_{m\to t}^{[i]})$ and $F_{t}^{[i]}=\mathbf{LN}(F_{t}^{[i]})$. Then, we put them into a cross-attention (CA) block:
@@ III-C Text-oriented Cross-Attention Module
Figure 3: The architecture of the Text-oriented Cross-attention module. The cross-attention block takes text-audio and text-video pairs as its input and the self-attention block takes text as its input. After the cross-attention block, two gates were introduced to eliminate adverse information. The memory gate decides how much proportion of the target modality’s components to be kept forwarding, and the fuse gate decides how much proportion of fused components to be injected into the target modality.
@@
III-D Unimodal Joint Learning
@@ IV Experiments
IV-A Experimental Settings
Datasets. We evaluate TCAN on CMU-MOSI [38] and CMU-MOSEI [39] datasets. To achieve a more comprehensive comparison, the experiments are conducted under word-unaligned settings. Table I depicts the composition of these two datasets. The CMU-MOSI dataset comprises 2,199 short monologue video clips. The acoustic features are extracted at a sampling rate of 12.5 Hz, while the visual features are extracted at a sampling rate of 15 Hz. The CMU-MOSEI dataset consists of a significantly larger number of samples compared to CMU-MOSI, with a total of 22,856 movie review video clips sourced from YouTube (approximately 10 times the size of CMU-MOSI). The acoustic features in CMU-MOSEI are extracted at a sampling rate of 20 Hz, while the visual features are extracted at a sampling rate of 15 Hz. For both CMU-MOSI and CMU-MOSEI datasets, each sample is annotated with a sentiment score ranging from -3 to 3. The sentiment scores represent different levels of sentiment, including highly negative, negative, weakly negative, neutral, weakly positive, positive, and highly positive. Following previous works [27, 10], we evaluate the model performance with the following metrics: the 7-class accuracy (Acc-7), the binary accuracy (Acc-2) where zero excluded evaluation results are used, the mean absolute error (MAE), the correlation of the model’s prediction with human (Corr), and the F1 score.
@@ IV-A Experimental Settings
TABLE I: The compositions of the number of samples in the CMU-MOSI and the CMU-MOSEI dataset.
@@
TABLE II: Comparison results on the CMU-MOSI. Bold is the best.
@@
TABLE III: Comparison results on the CMU-MOSEI dataset. Bold is the best.
@@
IV-B Comparisons with the State-of-the-Art
@@
TABLE IV: Ablation study on the influence of modality on the CMU-MOSEI dataset. The unimodal results are adapted from Mult [1].
@@
TABLE V: Ablation study on the influence of the center modality on the CMU-MOSI.
@@
TABLE VI: Ablation study on the influence of different network modules on the CMU-MOSI.
@@ IV-C Ablation Study
We separately remove each modality to investigate the performance of our proposed model. As illustrated in Table IV, we can see that the performance of “T only” is remarkably better than “A only” and “V only”, which to some extent implies the importance of text for MSA. When introducing text-oriented cross-attention modules, “TV pairs” or “TA pairs”, the performance has an apparent gain. Then, combining “TV pairs” and “TA pairs” further improves the recognition performance. In addition, the performance drop is relatively large when the visual modality is removed (comparing “TV pairs” and “TA pairs” related to “TV+TA pairs”).
Figure 4: Performance of TCAN with different hyperparameter $N$ on the CMU-MOSI. Here we use $F_{1}$ scores to show the performance.
To illustrate the dominance of the text modality, we substitute the other two modalities as the dominant modality for the experiments. As illustrated in Table V, the acoustic- and visual-oriented models consistently experience noticeable performance degradation. These observations indicate that the text modality possesses richer semantics and less noise, resulting in better feature reinforcement of the other two modalities.
Figure 5: Some representative examples for the visualization analysis of MSA. Red represents positive sentiment and blue represents negative sentiment.
We explore the importance of the proposed components by removing the gated mechanism and unimodal joint training separately. We can see that the gain degrades when removing one of the modules. As shown in Table VI, these observations suggest that gated mechanism and joint training modules play an important role in the result.
Finally, to investigate the impact of the number $N$ of the Text-oriented Cross-attention module on the model’s performance, we carried out experiments on the CMU-MOSI dataset using varying values for parameter $N$. The results are illustrated in Fig. 4. As $N$ increases, we observe a pattern in which the F1 scores initially rise and then decline, with the network achieving its optimal performance when $N=5$. We believed that the larger $N$ could result in better performance. Nevertheless, experiments indicate that an excessive number of layers may restrict the text modality’s capacity to guide the other two modalities. We should choose the appropriate network for different datasets, which is exactly what our proposed TCAN can flexibly do. When transitioning our model to a more intricate dataset, we can strategically augment the number of Text-Oriented Cross-modal Attention modules to attain optimal performance.
IV-D Visualization Analysis